A 91-day Treasury bill with a face value of $1 million is sold to yield 5.27 percent.

a. At what price did the T-bill sell if the yield was quoted by the market? (Use 365 days a year. Round the final answer to 2 decimal places.)

Price           $

b. At what price did the T-bill sell if the yield was an effective annual yield? (Use 365 days a year. Round the final answer to 2 decimal places.)

Price           $
Answer #1

1.
=(1-5.27%*91/365)*1000000=986861.0959

2.
=1000000/((1+5.27%)^(91/365))=987277.231355761